# EmoDubber: Towards High Quality and Emotion Controllable Movie Dubbing

### v2

@@ Abstract
Given a piece of text, a video clip, and a reference audio, the movie dubbing task aims to generate speech that aligns with the video while cloning the desired voice. The existing methods have two primary deficiencies: (1) They struggle to simultaneously hold audio-visual sync and achieve clear pronunciation; (2) They lack the capacity to express user-defined emotions. To address these problems, we propose EmoDubber, an emotion-controllable dubbing architecture that allows users to specify emotion type and emotional intensity while satisfying high-quality lip sync and pronunciation. Specifically, we first design Lip-related Prosody Aligning (LPA), which focuses on learning the inherent consistency between lip motion and prosody variation by duration level contrastive learning to incorporate reasonable alignment. Then, we design Pronunciation Enhancing (PE) strategy to fuse the video-level phoneme sequences by efficient conformer to improve speech intelligibility. Next, the speaker identity adapting module aims to decode acoustics prior and inject the speaker style embedding. After that, the proposed Flow-based User Emotion Controlling (FUEC) is used to synthesize waveform by flow matching prediction network conditioned on acoustics prior. In this process, the FUEC determines the gradient direction and guidance scale based on the user’s emotion instructions by the positive and negative guidance mechanism, which focuses on amplifying the desired emotion while suppressing others. Extensive experimental results on three benchmark datasets demonstrate favorable performance compared to several state-of-the-art methods.
@@ 2.1 Visual Voice Cloning
The V2C requires generating a waveform representing how a text might be said, but in step with the lip movements portrayed by a character, and in vocal style exemplified by reference audio. Some works focus on improving speaker identity to handle multi-speaker scenes [3, 7, 50, 28, 15]. For example, Speak2Dub [50] introduces speaker embedding extracted by pre-trained GE2E to phoneme encoder and mel-spectrogram decoder by learnable style affine transform, while StyleDubber [7] propose multi-scale style adaptor with phoneme and utterance level to strengthen speaker’s characteristics. Besides, some works attempt to combine visual representation to enhance prosody expressiver [16, 24, 6, 51]. For example, HPMDubbing [6] is a hierarchical dubbing method by bridging acoustic details with visual information: lip motion, face region, and scene. To improve the contextual prosody, MCDubber [51] enlarges the modeling object from single sentence to previous and following sentences, which incorporates more contextual video scenes. Although the speaker identity and prosody modeling have received attention, existing works still suffer from poor lip-sync and lifeless emotional expression, which is unacceptable in dubbing. In this work, we propose EmoDubber, a controllable emotional dubbing architecture to help users specify emotion they need in video, while bringing high quality lip-sync and pronunciation.
@@ 3.1 Overview
specifically, $E=\{c,\alpha,\beta\}$, where $c$, $\alpha$, and $\beta$ are emotion label, positive weight, and negative weight, respectively. The main architecture of the proposed model is shown in Figure 2. First, the Lip-related Prosody Aligning (LPA) module absorbs the input information (i.e., $T_{p}$, $R_{a}$, and $V_{l}$) to generate the lip-prosody context sequences with consistent duration cues, like basic pause and speech rate, which are guided by the proposed duration level contrastive learning. Next, the Pronunciation Enhancing (PE) strategy focuses on expanding phoneme sequences to video level by monotonic alignment search (MAS) and fusing it with the output of LPA by efficient conformer. Then, the Speaker Identity Adapting (SIA) module further decodes the output of the PE to acoustics prior information while introducing style from speaker identity. Finally, the Flow-based User Emotion Controlling (FUEC) renders emotions in the flow-matching vector field prediction process using Positive and Negative Guidance Mechanisms (PNGM) according to the user’s instructions $E$ to generate desired emotional speech. We detail each module below.
@@ 3.2 Lip-related Prosody Aligning
The proposed Lip-related Prosody Aligner (LPA) takes the script, silent video clips, and reference audio as input, and outputs the lip-prosody context sequences, which learn the pause and speed to align video by duration-level contrastive learning.
@@ 3.5 Flow-based User Emotion Controlling
where $v_{t}(\phi_{t}(x)|\mu,\theta)$ is the predicted gradient vector field of $\phi_{t}(x)$ according to the acoustics prior information $\mu$. Then, we can solve the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu,\theta)dt$ from $t=0$ to $t=1$ to generate the target mel-spectrogram $\hat{M}$ from noise $x_{0}$. Instead of Matcha-TTS [34], we carefully design two kinds of style affine learning in FMPN (intra-blocks and inter-blocks) to adapt multi-speaker scenarios effectively.
@@ 4.1 Implementation Details
Video frames are sampled at 25 FPS and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following [32, 29]. The window length, frame size, and hop length in STFT are 640, 1,024, and 160, respectively. The flow-prediction network is pre-trained on LibriSpeech [38] using 2 downsampling blocks, 2 midblocks, and 2 upsampling blocks. Each block had one Transformer layer with hidden dimensionality 256 and style affine layers [7], 2 heads, attention dimensionality 64, and snakebeta [25] activations. We use 8 heads for multi-head attention in LPA with 256 hidden sizes. The temperature coefficient $\tau$ of $\mathcal{L}_{cl}$ as 0.1. There are 5 conformer blocks in AVEC. We trained V2C, Chem, and GRID with batch sizes 64, 32, and 64, respectively. The Emotion Expert Classifier $\psi$ is trained on 13 emotional datasets with more than 50,000 emotional audio recordings in Emobox [31], which collects large-scale Speech Emotion Recognition (SER) benchmarks. During the inference process, $\gamma$ is set to 15, while $\alpha\in[0,5]$ and $\beta\in[0,2]$. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU. More details about the model and training are given in the supplementary material.
@@ 4.2 Datasets
V2C-Animation is a multi-speaker dataset for animation movie dubbing with identity and emotion annotations [3]. It is collected from 26 Disney cartoon movies and covers 153 diverse characters. The whole dataset has 10,217 video clips with paired audio and subtitles. The training/validation/test size are 60%, 10%, 30%.
@@ 4.3 Evaluation Metrics
Audio-visual Sync Evaluation. To evaluate the synchronization between the generated speech and the video quantitatively, we adopt Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) as our metrics, which are widely used to lip reading [49], talking face [48, 19], and video dubbing task [16, 28]. These metrics are based on the pre-trained SyncNet [5], which can explicitly test for lip synchronization in unconstrained videos in the wild [40, 5]. Compared to the length metric MCD-SL [3], we believe that LSE-C and LSE-D can more accurately measure the synchronization of vision and audio. The discussion of the two kinds of metrics is in Appendix E.
@@ 4.4 Comparison with SOTA Dubbing Methods
To compare with SOTA dubbing model without the function of emotion control, we remove the PNGM, i.e., maintaining Figure 2 (a)-(c) and flow matching prediction network to generate waveform. We compare with the recent dubbing baselines to comprehensively analyze. More details about baselines and V2C results are in Appendix.
@@ 4.5 Emotional Controlling Evaluation
Intensity Controlling Results. Figure 3 shows the intensity-controlling results of EmoDubber on the Chem benchmark. The higher Intensity Score means a stronger emotional intensity of speech. We vary $\alpha$ from 0.0 to 5.0 and $\beta$ from 0.0 to 2.0 and present results for two cases: one with a fixed $\beta$ value and another where $\beta$ varies with $\alpha$ to compare with the traditional baseline with no dual-scale guidance. Compared to the baseline, EmoDubber offers a wider range of intensity control, evident from our broader Intensity Score range from $\alpha=0.0$ to $\alpha=5.0$. Additionally, our method enables stronger emotional modulation, significantly achieving a higher Intensity Score than the baseline when $\alpha=5.0$. This demonstrates that our EmoDubber supports a broader range of emotional control and enables the generation of audio with stronger emotional intensity by FUEC. We also note that the combination values of $\alpha$ and $\beta$ can be more diverse, supporting a wider range of emotional intensity controls. More results about GRID and V2C are in Appendix B.
@@
Qualitative Results. To demonstrate the effect of EmoDubber’s emotion synthesis, we invite three volunteers to express their expected emotions and intensities on three videos from dubbing datasets, respectively. We visualize the mel-spectrograms and provide comparison using FUEC or not. For instance, in (a), the user selects “surprise” (c=4) with an intensity of $\alpha=3.1,\beta=0.3$, which results in a noticeable rise trend at the end of the spectrograms (see blue box). In (b), the user selects “angry” (c=6) with an intensity of $\alpha=4.5,\beta=0.9$, which brings a remarkable increase in energy. It indicates that a strong tone is being expressed to render angry. Finally, in (c), the user selects “sad” (c=2) with an intensity of $\alpha=5.0,\beta=1.7$. This relatively high-intensity setting produces a spectrum with diminished high-frequency energy and softened transitions, typical of frustrated expressions associated with sadness. Appendix A and D provide more visualization results.
@@ 5 Conclusion
In this work, we propose EmoDubber, a controllable emotion dubbing architecture to help users specify the emotion they need while satisfying high-quality lip sync and clear pronunciation. The lip-related prosody aligning learns the inherent consistency by duration-level contrastive learning to reason the correct audio-visual alignment. Building on contextual lip motion and prosody information, the proposed pronunciation enhancing strategy fuses the video-level phoneme sequence to improve intelligibility. Besides, the Flow-based User Emotion Controlling (FUEC) with positive and negative guidance dynamically adjusts the flow-matching prediction process to control emotion intensity flexibly. Extensive experiments on three widely adopted benchmarks show favorable performance.

### v3

@@ Abstract
Given a piece of text, a video clip, and a reference audio, the movie dubbing task aims to generate speech that aligns with the video while cloning the desired voice. The existing methods have two primary deficiencies: (1) They struggle to simultaneously hold audio-visual sync and achieve clear pronunciation; (2) They lack the capacity to express user-defined emotions. To address these problems, we propose EmoDubber, an emotion-controllable dubbing architecture that allows users to specify emotion type and emotional intensity while satisfying high-quality lip sync and pronunciation. Specifically, we first design Lip-related Prosody Aligning (LPA), which focuses on learning the inherent consistency between lip motion and prosody variation by duration level contrastive learning to incorporate reasonable alignment. Then, we design Pronunciation Enhancing (PE) strategy to fuse the video-level phoneme sequences by efficient conformer to improve speech intelligibility. Next, the speaker identity adapting module decodes acoustics prior and inject the speaker style embedding. After that, the proposed Flow-based User Emotion Controlling (FUEC) is used to synthesize waveform by flow matching prediction network conditioned on acoustics prior. In this process, the FUEC determines the gradient direction and guidance scale based on the user’s emotion instructions by the positive and negative guidance mechanism, which focuses on amplifying the desired emotion while suppressing others. Extensive experimental results demonstrate favorable performance compared to several state-of-the-art methods. The code and trained models will be made available at https://github.com/GalaxyCong/DubFlow.
@@ 2.1 Visual Voice Cloning
With the rapid development of deep learning [52, 53, 50, 62, 54, 31, 60, 51, 3, 10, 64] and speech synthesis [58, 12, 59], movie dubbing (i.e., V2C) has advanced in recent years. It needs to generate a waveform representing how a text might be said and emphasize to keep in step with the lip motion portrayed by a character. Some works focus on improving speaker identity to handle multi-speaker scenes [4, 8, 63, 32, 18]. For example, StyleDubber [8] propose multi-scale style adaptor with phoneme and utterance level to strengthen speaker’s characteristics. Besides, some works attempt to combine visual representation to enhance prosody expressiver [19, 27, 7, 66, 65, 67]. For example, HPMDubbing [7] is a hierarchical dubbing method by bridging acoustic details with visual information: lip motion, face region, and scene. Although the speaker identity and prosody modeling have received attention, existing works still suffer from poor lip-sync and lifeless emotional expression, which is unacceptable in dubbing. In this work, we propose EmoDubber, a controllable emotional dubbing architecture to help users specify emotion they need in video, while bringing high quality lip-sync and pronunciation.
@@ 3.1 Overview
specifically, $E=\{c,\alpha,\beta\}$, where $c$, $\alpha$, and $\beta$ are emotion label, positive weight, and negative weight, respectively. The main architecture of the proposed model is shown in Figure 2. First, the Lip-related Prosody Aligning (LPA) module absorbs the input information (i.e., $T_{p}$, $R_{a}$, and $V_{l}$) to generate the lip-prosody context sequences with consistent duration cues, like basic pause and speech rate, which are guided by the proposed duration level contrastive learning. Next, the Pronunciation Enhancing (PE) strategy focuses on expanding phoneme sequences to video level by monotonic alignment search (MAS) and fusing it with the output of LPA by efficient conformer. Then, the Speaker Identity Adapting (SIA) module further decodes the output of the PE to acoustics prior information while introducing style from speaker identity. Finally, the Flow-based User Emotion Controlling (FUEC) renders emotions in the flow-matching vector field prediction process using Positive and Negative Guidance Mechanisms (PNGM) according to the user’s instructions $E$ to generate desired emotional speech.
@@ 3.2 Lip-related Prosody Aligning
The core of Lip-related Prosody Aligning (LPA) is duration-level contrastive learning (DLCL), which aims to capture the correct relevance between prosody and lip motion to bring implicit modeling while focusing on supplying reasonable explicit duration modeling for pronunciation enhancing (PE) strategy.
@@ 3.5 Flow-based User Emotion Controlling
where $v_{t}(\phi_{t}(x)|\mu,\theta)$ is the predicted gradient vector field of $\phi_{t}(x)$ according to the acoustics prior information $\mu$. Then, we can solve the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu,\theta)dt$ from $t=0$ to $t=1$ to generate the target mel-spectrogram $\hat{M}$ from noise $x_{0}$. Besides, we use the style affine learning in FMPN to adapt multi-speaker scenarios effectively.
@@ 4.1 Implementation Details
Video frames are sampled at 25 FPS and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following [36, 33]. The window length, frame size, and hop length in STFT are 640, 1,024, and 160, respectively. Each block had one Transformer layer with hidden dimensionality 256 and style affine layers [8], 2 heads, attention dimensionality 64, and snakebeta [28] activations. We use 8 heads for multi-head attention in LPA with 256 hidden sizes. The temperature coefficient $\tau$ of $\mathcal{L}_{cl}$ as 0.1. There are 5 conformer blocks in AVEC. We trained V2C, Chem, and GRID with batch sizes 64, 32, and 64, respectively. The Emotion Expert Classifier $\psi$ is trained on 13 emotional datasets with more than 50,000 emotional audio recordings in Emobox [35], which collects large-scale Speech Emotion Recognition (SER) benchmarks. The structure of the classifier is a 6-layer 1D CNN with BatchNorm, LeakyReLU, and Dropout. Only phoneme encoder, USL, and flow decoder are pre-trained on LibriSpeech [42] to fair compare with the pre-trained dub method [63]. During the inference process, $\gamma$ is set to 15, while $\alpha\in[0,9]$ and $\beta\in[0,2]$. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU. More details are given in the Appendix.
@@ 4.3 Evaluation Metrics
Audio-visual Sync Evaluation. To evaluate the synchronization between the generated speech and the video quantitatively, we adopt Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) as our metrics, which are widely used to lip reading [61], talking face [57, 22], and video dubbing task [19, 32]. These metrics are based on the pre-trained SyncNet [6], which can explicitly test for lip synchronization in unconstrained videos in the wild [44, 6]. Compared to the length metric MCD-SL [4], we believe that LSE-C and LSE-D can more accurately measure the synchronization of vision and audio.
@@ 4.4 Comparison with SOTA Dubbing Methods
To compare with SOTA dubbing model without the function of emotion control, we remove the PNGM, i.e., maintaining Figure 2 (a)-(c) and flow matching prediction network to generate waveform. We compare with the recent dubbing baselines to comprehensively analyze.
@@ 4.5 Emotional Controlling Evaluation
Intensity Controlling Results. Figure 3 shows the intensity-controlling results of EmoDubber on the Chem benchmark. The higher Intensity Score means a stronger emotional intensity of speech. We vary $\alpha$ from 0.0 to 5.0 and $\beta$ from 0.0 to 2.0 and present results for two cases: one with a fixed $\beta$ value and another where $\beta$ varies with $\alpha$ to compare with the traditional baseline with no dual-scale guidance. Compared to the baseline, EmoDubber offers a wider range of intensity control, evident from our broader Intensity Score range from $\alpha=0.0$ to $\alpha=5.0$. Additionally, our method enables stronger emotional modulation, significantly achieving a higher Intensity Score than the baseline when $\alpha=5.0$. This demonstrates that our EmoDubber supports a broader range of emotional control and enables the generation of audio with stronger emotional intensity by FUEC. We also note that the combination values of $\alpha$ and $\beta$ can be more diverse, supporting a wider range of emotional intensity controls.
@@
Qualitative Results. To demonstrate the effect of EmoDubber’s emotion synthesis, we invite three volunteers to express their expected emotions and intensities on three videos from dubbing datasets, respectively. We visualize the mel-spectrograms and provide comparison using FUEC or not. For instance, in (a), the user selects “surprise” (c=4) with an intensity of $\alpha=3.1,\beta=0.3$, which results in a noticeable rise trend at the end of the spectrograms (see blue box). In (b), the user selects “angry” (c=6) with an intensity of $\alpha=4.5,\beta=0.9$, which brings a remarkable increase in energy. It indicates that a strong tone is being expressed to render angry. Finally, in (c), the user selects “sad” (c=2) with an intensity of $\alpha=5.0,\beta=1.7$. This relatively high-intensity setting produces a spectrum with diminished high-frequency energy and softened transitions, typical of frustrated expressions associated with sadness.
@@ 5 Conclusion
In this work, we propose EmoDubber, a controllable emotion dubbing architecture to help users specify the emotion they need while satisfying high-quality lip sync and clear pronunciation. The LPA learns the inherent consistency by duration-level contrastive learning to reason the correct audio-visual alignment. Building on contextual information, the proposed PE strategy fuses the explicit phoneme chunk on video level to improve pronunciation by efficient conformer. Besides, the Flow-based User Emotion Controlling (FUEC) with positive and negative guidance dynamically adjusts the flow-matching prediction process to control emotion intensity flexibly. Extensive experiments on real dubbing benchmarks show favorable performance.
6 Acknowledgement
This work was supported by National Natural Science Foundation of China: 62322211, 61925201, 62132001, 62432001, and Beijing Natural Science Foundation (L247006), “Pioneer” and “Leading Goose” R&D Program of Zhejiang Province (2024C01023), Key Laboratory of Intelligent Processing Technology for Digital Music (Zhejiang Conservatory of Music), Ministry of Culture and Tourism (2023DMKLB004). Yuankai Qi, Anton van den Hengel, and Jian Yang are not supported by the aforementioned fundings.